# CIVIL: Causal and Intuitive Visual Imitation Learning

### v2

@@ III-A Using Low-Dimensional Features to Accelerate Learning
When the error follows a normal distribution, the ordinary least squares estimator $\hat{\theta}$ is also normally distributed [46]:
@@
The variance in the estimated parameters corresponds to the uncertainty in converging to the human’s policy. We quantify this uncertainty using the continuous Shannon entropy of the estimator’s distribution [47]:
$h(f_{\hat{\theta}})=\frac{d}{2}+\frac{d}{2}\ln 2\pi+\frac{1}{2}\ln|\Sigma_{\hat{\theta}}|$
@@
From the definition of variance [48], the expected value of the outer product can be written in terms of the mean and variance of the feature vectors:
@@
To simplify this further, we assume that the feature vectors are normally distributed such that $\phi_{i}\sim(0,\sigma_{\phi}^{2}I_{d})$. Note that this is a common assumption in previous approaches that use Variational Autoencoders (VAEs) [49] to encode image data. Equipped with this assumption, we can now write the expected value of the outer product as $\mathbb{E}[\Phi^{T}\Phi]=N\sigma_{\phi}^{2}I_{d}$ and substitute it back into the covariance of the estimator distribution:
@@
Proposition 1 illustrates the importance of mapping the robot’s high-dimensional observations into a minimal feature representation. But what is the right representation to use? Thus far we have assumed that the robot has access to a feature function $\psi$ that extracts sufficient information for learning the task. In the next subsection, we will show that there can be many feature representations that are sufficient for imitating the actions in the training data but do not align with the human’s reasoning $\phi^{*}$. As a result, these alternate feature representations are susceptible to covariate shift and fail when the robot encounters new states at test time.
III-B Causal Confusion in Visual Imitation Learning
@@
More generally, these correlations create causal confusion [1] when applying the learned feature function in unseen scenarios. To demonstrate this formally, we return to the linear regression problem. We now assume that the encoder matrix $\psi$ is unknown and combine it with the policy parameters to create a single weight matrix $W=\psi\theta$:
@@
Here $[XY]^{{\dagger}}$ denotes the pseudo-inverse of $[XY]$. While there are infinite ways to factorize $\hat{W}$ into the individual components $\hat{\psi}$ and $\hat{\theta}$, because we have a unique $\hat{W}$, all of these choices are equivalent to the human’s true weights $W^{*}$:
@@
This difference in learned weights will not affect the robot’s performance if the correlations in the training data are also present at test time. In this case, the test inputs $[XY]_{test}$ will have the same null space as the training data:
@@
As a result, $\hat{W}$ will produce the same actions as $W$. However, if the correlations in the inputs change at test time, then the learned weights $\hat{W}$ will not produce the same actions as $W^{*}$ for any non-trivial $V$. For instance, if the positions of the bowl and cup are no longer related during testing, the test inputs $[X,Y]_{test}$ will have full rank. This means that $V$ will not belong to the null space of $[X,Y]_{test}$:
@@ IV Causal and Intuitive
In Section IV-A we first equip humans with instruments that enable them to intuitively communicate information about the relevant features (i.e., $\phi^{*}$) and which parts of the observations they consider when extracting these features (i.e, $\psi^{*}$). Next, in Section IV-B, we describe our network architecture for extracting features from robot observations and mapping them to corresponding robot actions. We apply this architecture in Section IV-C to develop the CIVIL algorithm which synthesizes data collected from our instruments to align the robot’s features with the human’s reasoning. Finally, in Section IV-D we provide implementation details. A key contribution of our approach is that the robot does not need instruments after training and can perform the task autonomously at test time based only on visual observations.
@@
IV-A Obtaining Task-Relevant Information from Humans
@@
Communicating Relevant Features. Task actions often depend on contextual variables such as the position of a target object, the color of a traffic signal, or the speed of a moving obstacle. We can enable humans to communicate these variables to the robot by equipping them with the required sensors and interfaces. In this work, we provide humans with physical markers to specify poses and waypoints relevant to the desired task. Specifically, we let humans place ArUco markers [50] in the environment before providing demonstrations. These markers then continuously stream their poses $b\in\mathbb{R}^{d_{b}}$ to the robot as the human performs the task. The ArUco markers have a binary pattern that can be detected by the robot’s camera for pose estimation; these markers are also small (one inch in width), lightweight (${\sim}10$ grams), and adhere to various surfaces in the environment. Consider our running example in Figure 2: when teaching the robot to pick a coffee cup, humans may attach a marker to its side to indicate where they want to grasp. The marker poses $b$ directly inform the task actions (i.e. how the human teleoperates the robot). Hence, we consider poses $b$ as task-relevant features that the robot learner should extract from its observations.
@@
To connect the human’s utterances with visual observations we leverage a language-conditioned video segmentation model, DEVA [51]. In practice, DEVA associates the human’s verbal prompts with objects in the robot’s view, producing bounding boxes ${}\{\mathcal{B}\}$ around objects the human mentions. Note that humans can also indicate relevant objects using ArUco markers. We therefore harness the markers similarly to language, and give them a dual purpose: in addition to estimating marker pose, we detect objects closest to the marker and retrieve those objects’ bounding boxes. Similar to the human teacher, the robot should focus on the visual elements within the bounding boxes when extracting the features. We expect that this attention will reduce causal confusion with irrelevant objects and allow the robot to implicitly infer task-relevant features from the demonstration data.
Data Collection. Overall, we shift the demonstration process so that human teachers can use physical markers to explicitly convey relevant poses, and natural language (or markers) to indicate relevant objects for implicit features. Figure 2 shows how we integrate these instruments into the learning pipeline. We ask humans to place the markers before providing demonstrations (Step 1), and then issue natural language commands as they demonstrate the task (Step 2). Our experimental data suggests that both instruments are intuitive for humans to deploy. In our studies, users required less than $15$ seconds to attach the markers to relevant objects for teaching a coffee-making task (see Section VII). The robot stores the $(b,l)$ data collected from these instruments alongside the states $x$, images $y$, and actions $u$. Thus, the augmented dataset $\mathcal{D}$ contains information about what actions to imitate and why in the form of $(x,y,u,b,l)$ tuples. Once all demonstrations have been collected, humans remove the markers from the environment (Step 3). We inpaint these markers from images in the dataset so that the robot does not need to rely on seeing the markers in order to perform the task. The robot only retains the marker poses it recorded during the offline demonstrations.
@@
IV-B Network Architecture
Our proposed network architecture is illustrated in Figure 3. The robot uses an encoder network $f_{\psi}(x,y)=\phi$ to map the $n$-dimensional visual observations $y\in\mathbb{R}^{n}$ into $d$-dimensional features $\phi\in\mathbb{R}^{d}$. Based on Proposition 1 in Section III-A, designers should set the dimensionality of these features to be much lower than that of the observations (i.e., $d<<n$) in order to accelerate robot learning. As we will describe later, $d$ also depends on the number of markers or language utterances that the human teacher provides.
@@ IV-B Network Architecture
Aspects of our architecture follow the structure of previous visual imitation learning approaches [52, 53]. But — as we will show — the key difference is how we employ supplementary inputs $(b,l)$ to align the learned features with the human’s true features. In what follows we introduce the auxiliary networks and losses needed to achieve this alignment.
IV-C Supervised Learning with CIVIL
@@
Implicit features. Other than explicitly specifying the relevant positions through markers, the human also indicates relevant objects using natural language prompts $l\in\mathcal{L}$. Here we explain how the robot maps these prompts to the implicit features $\phi_{implicit}$ from Equation (8). Our first step is to locate the objects mentioned by the human within the corresponding image $y$. We do this by feeding the image $y$ and language prompt $l$ to a DEVA model to obtain a bounding box $\mathcal{B}$ for each mentioned object. We also generate bounding boxes for objects that overlap with any markers detected in the image. In this way, for each $(y,l)$ pair we obtain a set ${}\{\mathcal{B}\}$ that includes bounding boxes of task-relevant objects.
We recognize that the human teacher understands the desired task, and we assume that we can rely on that human to identify the key objects or aspects of the task via language and markers. This implies that parts of the image $y$ outside the bounding boxes ${}\{\mathcal{B}\}$ are likely irrelevant and should be ignored by the robot when extracting features. To enforce this, we generate masked images $y^{\prime}\in\mathbb{R}^{n}$ by setting all pixels in $y$ that are not within the bounding boxes as zero, and then incorporate these filtered images into the training dataset $\mathcal{D}$.
@@
$\mathcal{L}_{causal}=\sum_{(x,y,y^{\prime})\in\mathcal{D}}$ $||f_{\psi}(x,y^{\prime})-c_{\lambda}(x,y)||^{2}$ (14)
@@
CIVIL Algorithm. The steps for training our architecture are listed in Algorithm 1 (and visualized in Figure 3). The human first places markers in the environment to stream relevant poses and then demonstrates the task while providing natural language prompts to indicate the relevant objects. The robot uses the markers and language instructions to obtain bounding boxes for all key objects and mask the irrelevant portions of the robot images. These masked images are added to the training dataset along with the marker readings. We then train our network architecture end-to-end by minimizing the combined loss $\mathcal{L}_{civil}$ in the first training phase:
@@
IV-D Implementation
A public CIVIL repository is available here: https://github.com/CIVIL2025/Implementation.
During our experiments the robot takes images from both a static third-person view $y_{static}$ and an ego-centric view $y_{ego}$. Accordingly, we train different encoder networks $f_{\psi_{1}}(x,y_{static})=\phi_{static}$ and $f_{\psi_{2}}(x,y_{ego})=\phi_{ego}$ for images from each camera view, and implement two corresponding causal networks $c_{\lambda_{1}}$ and $c_{\lambda_{2}}$. While $\phi_{static}$ has both explicitly and implicitly learned components as in Equation (8), we only extract implicit features from the ego-centric view because it does not always observe the marker positions (i.e., objects move in and out of the ego frame). We pass both features $(x,\phi_{static},\phi_{ego})$ as input to the robot policy. Before feeding these inputs to the policy transformer $\pi_{\theta}$, we project the states and features into separate tokens of size $128$ and add a sinusoidal positional encoding to each token to indicate its location in the sequence [54]. The encoders are convolutional neural networks; specifically, we choose ResNet-18 initialized with pre-trained weights [55]. The policy architecture includes a 2-layer transformer encoder followed by a multi-layer perceptron (MLP) action network with two hidden layers.
@@ V Simulations
  • Behavior cloning (BC) [56]: A standard imitation learning approach. BC learns to encode camera images and map them to robot actions by only training the policy based on the human’s demonstrated actions. This approach forces the robot to implicitly infer the task-relevant features.
@@
  • Self-Supervised Features (BYOL) [17]: A self-supervised framework that learns image representations by mapping different views to the same feature encoding. The alternative views are generated using transformations such as random cropping, flipping, and color jittering. BYOL learns general visual features that are not supervised to align with the human’s intention. In our experiments we pre-trained a BYOL encoder on the images in the training data as well as the play data, froze it, and then used its self-supervised features to train the downstream policy.
@@
  • Object-Oriented Features (VIOLA) [4]: An approach that encodes images by focusing on objects in the scene. VIOLA uses a pre-trained Region Proposal Network (RPN) [57] to obtain bounding boxes for $k$ observed objects and then extracts object-specific features. In our simulations we provided VIOLA with perfect detection by giving it the ground-truth bounding boxes of all objects in the environment. We then randomly selected $k=5$ of these objects to extract object features as in the original implementation. We ensure that these objects include the task-relevant item. By segmenting known objects, VIOLA learns to ignore background variations. However, the robot still needs to figure out which of the $k$ objects are relevant by training the features and downstream policy to imitate human actions in an end-to-end manner. Note that — unlike our approach — VIOLA requires access to the object bounding boxes even during testing.
@@
  • Task-Specific Object Features (Task-VIOLA) [45]: This approach is a variation of VIOLA that enables human teachers to indicate the desired objects by scribbling on the robot’s images. The robot then obtains point clouds of the annotated objects from its depth camera, and extracts features by training the downstream policy to imitate human actions. We replace object point clouds with image segmentations for a fair comparison with other methods that only use RGB images. Note that (similar to VIOLA) this approach also requires a pre-trained vision-language model during test time to segment the objects.
  • Vision-Language Features (CLIP) [7]: The methods discussed so far only derive features from visual inputs. We now include a baseline that learns from both images and language prompts. Specifically, we use a CLIP encoder that associates visual concepts with their text descriptions by mapping both inputs to the same feature space. CLIP features trained on large text-image datasets are general-purpose and may not directly apply to downstream robot tasks [12]. Therefore, we take a pre-trained ResNet-based CLIP encoder RN50$\times$4 and fine-tune it for our simulation tasks by adding top and bottom adapter layers and then training them with the robot policy as in [5].
@@
Unlike these baselines our CIVIL approach leverages human inputs to supervise a causal feature embedding. In contrast to Task-VIOLA — which lets humans mark relevant objects on a computer screen — CIVIL does not need a pre-trained model to segment the objects at run time.
@@
Simulation Environment. We trained and evaluated all methods on three tasks within the CALVIN environment  [8] shown in Figure 4. This $3$D environment includes a $7$-DOF Franka Emika Panda robot arm, three differently colored cubes on a workbench, a sliding door, a drawer, a light bulb operated with a control switch, and an LED controlled with a button. We randomly initialize these elements during data collection and evaluation. The demonstrations are either simulated using a pre-trained expert policy [58] or manually collected by an expert teacher. We also had the human expert specify the relevant objects and obtained ground-truth poses of these objects from the simulation environment.
@@
Figure 5: Section V results for visual imitation learning in the simulated CALVIN environment. We compare our proposed approach (CIVIL) to standard behavior cloning (BC) and baselines that use self-supervised features (BYOL), object-specific features (VIOLA and Task-VIOLA), or vision-language features (CLIP). In the Picking task we observe that CIVIL significantly outperforms all baselines in picking up the red block from different positions across the table. Our approach is particularly effective for block positions that are outside the training distribution (i.e., center of the table). This is likely because CIVIL understands what aspects of the image should influence its policy, making the system robust to background clutter or shifting positions. In the Sliding task, we find that CIVIL successfully learns to move the drawer or slider based on the state of the light bulb in just $40$ demonstrations. In contrast, baselines that use pre-trained features (BYOL and CLIP) are less precise in detecting the light signal, which reduces their success rate. This suggests that CIVIL can also learn to extract non-positional features (e.g., light bulb state) more efficiently by masking its images based on human-provided language prompts. Lastly, in the Stacking task CIVIL leverages both markers and language to extract the positions of the pink and blue blocks as well as the state of the light bulb. This enables the robot to stack the red block more successfully on either the pink or blue block, resulting in a significantly higher success rate.
@@
Demonstrations. At each timestep of a task demonstration we record an RGB image $y_{env}$ of size $200\times 200$ from a static camera that observes the entire manipulation environment, an egocentric RGB image $y_{ego}$ of size $84\times 84$ from a gripper-mounted camera, an $8$-dimensional robot state $x$, and a $7$-dimensional end-effector action $u$. The state includes $7$ joint angles of the robot arm and a Boolean gripper state. The action is a $6$-dimensional linear and angular velocity and a Boolean gripper actuation. Additionally, we obtain bounding boxes ${}\{\mathcal{B}\}$ for all objects in the simulation environment and explicit features $b$ in the form of $6$-dimensional Cartesian poses of relevant objects in that task. Each method uses a different combination of these inputs to train the feature encoders and robot policy. For the Picking and Stacking tasks, we collect an equal amount of play data which includes images, bounding boxes, and relevant poses in randomly initialized scenarios, but it does not include robot states and expert actions.
@@
Results. Our results are summarized in Figure 5. Each method is trained on datasets having $10$, $20$, $40$, $80$, and $120$ demonstrations. For statistical robustness, we conduct $10$ independent training and testing runs for each method and dataset size. In each run, we test the trained policy across $100$ randomized task configurations and report the average success rate.
@@
Surprisingly, we found that standard behavior cloning performed well in this task. We attribute this result to the small size of its feature space. While BC only extracts one feature token for each camera, the object-centric methods extract two tokens: global and object-specific features. Following our analysis in Section III-A, a more compact feature space could enable BC to learn more efficiently. Overall, this simulation result illustrates that by masking images based on language prompts CIVIL is able to extract non-positional features that help it perform the task more successfully.

### v3

@@ 3.1 Using Low-Dimensional Features to Accelerate Learning
When the error follows a normal distribution, the ordinary least squares estimator $\hat{\theta}$ is also normally distributed amemiya1985advanced :
@@
The variance in the estimated parameters corresponds to the uncertainty in converging to the human’s policy. We quantify this uncertainty using the continuous Shannon entropy of the estimator’s distribution ahmed1989entropy :
[…]
@@
From the definition of variance casella2024statistical , the expected value of the outer product can be written in terms of the mean and variance of the feature vectors:
@@
To simplify this further, we assume that the feature vectors are normally distributed such that $\phi_{i}\sim(0,\sigma_{\phi}^{2}I_{d})$. Note that this is a common assumption in previous approaches that use Variational Autoencoders (VAEs) pinheiro2021variational to encode image data. Equipped with this assumption, we can now write the expected value of the outer product as $\mathbb{E}[\Phi^{T}\Phi]=N\sigma_{\phi}^{2}I_{d}$ and substitute it back into the covariance of the estimator distribution:
@@
Proposition 1 illustrates the importance of mapping the robot’s high-dimensional observations into a minimal feature representation. But what is the right representation? Thus far we have assumed that the robot has access to a feature function $\psi$ that extracts sufficient information for learning the task. In the next subsection, we will show that there can be many feature representations that are sufficient for imitating the actions in the training data but do not align with the human’s reasoning $\phi^{*}$. As a result, these alternate feature representations are susceptible to covariate shift and fail when the robot encounters new states at test time.
3.2 Causal Confusion in Visual Imitation Learning
@@
More generally, these correlations create causal confusion de2019causal when applying the learned feature function in unseen scenarios. To demonstrate this formally, we return to the linear regression problem. We now assume that the encoder matrix $\psi$ is unknown and combine it with the policy parameters to create a single weight matrix $W=\psi\theta$:
@@
Here $XY]^{{\dagger}}$ denotes the pseudo-inverse of $[XY]$. While there are infinite ways to factorize $\hat{W}$ into the components $\hat{\psi}$ and $\hat{\theta}$, because we have a unique $\hat{W}$, all of these choices are equivalent to the human’s true weights $W^{*}$:
@@
This difference in learned weights will not affect the robot’s performance if the correlations in the training data are also present at test time. In this case, the test inputs $XY]_{test}$ will have the same null space as the training data:
@@
As a result, $\hat{W}$ will produce the same actions as $W$. However, if the correlations in the inputs change at test time, then the learned weights $\hat{W}$ will not produce the same actions as $W^{*}$ for any non-trivial $V$. For instance, if the positions of the bowl and cup are no longer related during testing, the test inputs $X,Y]_{test}$ will have full rank. This means that $V$ will not belong to the null space of $X,Y]_{test}$:
@@ 4 Causal and Intuitive
In Section 4.1 we first equip humans with instruments that enable them to intuitively communicate information about the relevant features (i.e., $\phi^{*}$) and which parts of the observations they consider when extracting these features (i.e, $\psi^{*}$). Next, in Section 4.2, we describe our network architecture for extracting features from robot observations and mapping them to corresponding robot actions. We apply this architecture in Section 4.3 to develop the CIVIL algorithm which synthesizes data collected from our instruments to align the robot’s features with the human’s reasoning. Finally, in Section 4.4 we provide implementation details. A key contribution of our approach is that the robot does not need instruments after training and can perform the task autonomously at test time based only on visual observations.
@@
4.1 Obtaining Task-Relevant Information from Humans
@@
Communicating Relevant Features. Task actions often depend on contextual variables such as the position of a target object, the color of a traffic signal, or the speed of a moving obstacle. We can enable humans to communicate these variables to the robot by equipping them with the required sensors and interfaces. In this work, we provide humans with physical markers to specify poses and waypoints relevant to the desired task. Specifically, we let humans place ArUco markers garrido2014automatic in the environment before providing demonstrations. These markers then continuously stream their poses $b\in\mathbb{R}^{d_{b}}$ to the robot as the human performs the task. The ArUco markers have a binary pattern that can be detected by the robot’s camera for pose estimation; these markers are also small (one inch in width), lightweight (${\sim}10$ grams), and adhere to various surfaces in the environment. Consider our running example in Figure 2: when teaching the robot to pick a coffee cup, humans may attach a marker to its side to indicate where they want to grasp. The marker poses $b$ directly inform the task actions (i.e. how the human teleoperates the robot). Hence, we consider poses $b$ as task-relevant features that the robot learner should extract from its observations and incorporate in its control policy.
@@
To connect the human’s utterances with visual observations we leverage a language-conditioned video segmentation model, DEVA cheng2023tracking . In practice, DEVA associates the human’s verbal prompts with objects in the robot’s view, producing bounding boxes $\{\mathcal{B}\}$ around objects the human mentions. Note that humans can also indicate relevant objects using ArUco markers. We therefore harness the markers similarly to language, and give them a dual purpose: in addition to estimating marker pose, we detect objects closest to the marker and retrieve those objects’ bounding boxes. Similar to the human teacher, the robot should focus on the visual elements within the bounding boxes when extracting the features. We expect that this attention will reduce causal confusion with irrelevant objects and allow the robot to implicitly infer task-relevant features from the demonstration data.
Data Collection. Overall, we shift the demonstration process so that human teachers can use physical markers to explicitly convey relevant poses, and natural language (or markers) to indicate relevant objects for implicit features. Figure 2 shows how we integrate these instruments into the learning pipeline. We ask humans to place the markers before providing demonstrations (Step 1), and then issue natural language commands as they demonstrate the task (Step 2). Our experimental data suggests that both instruments are intuitive for humans to deploy. In our studies, users required less than $15$ seconds to attach the markers to relevant objects for teaching a coffee-making task (see Section 7). The robot stores the $(b,l)$ data collected from these instruments alongside the states $x$, images $y$, and actions $u$. Thus, the augmented dataset $\mathcal{D}$ contains information about what actions to imitate and why in the form of $(x,y,u,b,l)$ tuples. Once all demonstrations have been collected, humans remove the markers from the environment (Step 3). We inpaint these markers from images in the dataset so that the robot does not need to rely on seeing the markers in order to perform the task. The robot only retains the marker poses it recorded during the offline demonstrations.
@@
4.2 Network Architecture
Our proposed network architecture is illustrated in Figure 3. The robot uses an encoder network $f_{\psi}(x,y)=\phi$ to map the $n$-dimensional visual observations $y\in\mathbb{R}^{n}$ into $d$-dimensional features $\phi\in\mathbb{R}^{d}$. Based on Proposition 1 in Section 3.1, designers should set the dimensionality of these features to be much lower than that of the observations (i.e., $d<<n$) in order to accelerate robot learning. As we will describe later, $d$ also depends on the number of markers or language utterances that the human teacher provides.
@@ 4.2 Network Architecture
Aspects of our architecture follow the structure of previous visual imitation learning approaches brohan2022rt ; haldarbaku . But — as we will show — the key difference is how we employ supplementary inputs $(b,l)$ to align the learned features with the human’s true features. In what follows we introduce the auxiliary networks and losses needed to achieve this alignment.
4.3 Supervised Learning with CIVIL
@@
Implicit features. Other than explicitly specifying the relevant positions through markers, the human also indicates relevant objects using natural language prompts $l\in\mathcal{L}$. Here we explain how the robot maps these prompts to the implicit features $\phi_{implicit}$ from Equation (8). Our first step is to locate the objects mentioned by the human within the corresponding image $y$. We do this by feeding the image $y$ and language prompt $l$ to a DEVA model to obtain a bounding box $\mathcal{B}$ for each mentioned object. We also generate bounding boxes for objects that overlap with any markers detected in the image. For each $(y,l)$ pair we thus obtain a set $\{\mathcal{B}\}$ that includes bounding boxes of task-relevant objects.
We recognize that the human teacher understands the desired task, and we assume that we can rely on that human to identify the key objects or aspects of the task via language and markers. This implies that parts of the image $y$ outside the bounding boxes $\{\mathcal{B}\}$ are likely irrelevant and should be ignored by the robot when extracting features. To enforce this, we generate masked images $y^{\prime}\in\mathbb{R}^{n}$ by setting all pixels in $y$ that are not within the bounding boxes as zero, and incorporate these filtered images into the training dataset $\mathcal{D}$.
@@
[…]
@@
CIVIL Algorithm. The steps for training our architecture are listed in Algorithm 1 (and visualized in Figure 3). The human first places markers in the environment to stream relevant poses and then demonstrates the task while providing natural language prompts to indicate the relevant objects. The robot uses the markers and language instructions to obtain bounding boxes for all key objects and mask the irrelevant portions of the robot images. These masked images are added to the training dataset along with the marker readings. We train our network architecture end-to-end by minimizing the loss $\mathcal{L}_{civil}$ in the first training phase:
@@
4.4 Implementation
A public CIVIL repository q available here: https://github.com/CIVIL2025/Implementation.
During our experiments the robot takes images from both a static third-person view $y_{static}$ and an ego-centric view $y_{ego}$. Accordingly, we train different encoder networks $f_{\psi_{1}}(x,y_{static})=\phi_{static}$ and $f_{\psi_{2}}(x,y_{ego})=\phi_{ego}$ for images from each camera view, and implement two corresponding causal networks $c_{\lambda_{1}}$ and $c_{\lambda_{2}}$. While $\phi_{static}$ has both explicitly and implicitly learned components as in Equation (8), we only extract implicit features from the ego-centric view because it does not always observe the marker positions (i.e., objects move in and out of the ego frame). We pass both features $(x,\phi_{static},\phi_{ego})$ as input to the robot policy. Before feeding these inputs to the policy transformer $\pi_{\theta}$, we project the states and features into separate tokens of size $128$ and add a sinusoidal positional encoding to each token to indicate its location in the sequence vaswani2017attention . The encoders are convolutional neural networks; specifically, we choose ResNet-18 initialized with pre-trained weights he2016deep . The policy architecture includes a 2-layer transformer encoder followed by a multi-layer perceptron (MLP) action network with two hidden layers.
@@ 5 Simulations
  • Behavior cloning (BC) jain2019learning : A standard imitation learning approach. BC learns to encode camera images and map them to robot actions by only training the policy based on the human’s demonstrated actions. This approach forces the robot to implicitly infer the task-relevant features.
@@
  • Self-Supervised Features (BYOL) grill2020bootstrap : A self-supervised framework that learns image representations by mapping different views to the same feature encoding. The alternative views are generated using transformations such as random cropping, flipping, and color jittering. BYOL learns visual features that are not supervised to align with the human’s intention. In our experiments we pre-trained a BYOL encoder on the images in the training data as well as the play data, froze it, and then used its self-supervised features to train the downstream policy.
@@
  • Object-Oriented Features (VIOLA) zhu2023viola : An approach that encodes images by focusing on objects in the scene. VIOLA uses a pre-trained Region Proposal Network (RPN) ren2015faster to obtain bounding boxes for $k$ observed objects and then extracts object-specific features. In our simulations we provided VIOLA with perfect detection by giving it the ground-truth bounding boxes of all objects in the environment. We then randomly selected $k=5$ of these objects to extract object features as in the original implementation. We ensure that these objects include the task-relevant item. By segmenting known objects, VIOLA learns to ignore background variations. However, the robot still needs to figure out which of the $k$ objects are relevant by training the features and downstream policy to imitate human actions in an end-to-end manner. Note that — unlike our approach — VIOLA requires access to the object bounding boxes even during testing.
@@
  • Task-Specific Object Features (Task-VIOLA) zhu2023learning : This approach is a variation of VIOLA that enables human teachers to indicate the desired objects by scribbling on the robot’s images. The robot then obtains point clouds of the annotated objects from its depth camera, and extracts features by training the downstream policy to imitate human actions. We replace object point clouds with image segmentations for a fair comparison with other methods that only use RGB images. Note that (similar to VIOLA) this approach also requires a pre-trained vision-language model during test time to segment the objects.
[…]
@@
Unlike these baselines our CIVIL approach leverages human inputs from the demonstration process to supervise a causal feature embedding. In contrast to Task-VIOLA — which lets humans mark relevant objects on a computer screen — CIVIL does not need a pre-trained model to segment the objects at run time.
@@
Simulation Environment. We trained and evaluated all methods on three tasks within the CALVIN environment  mees2022calvin shown in Figure 4. This 33D environment includes a $7$-DOF Franka Emika Panda robot arm, three differently colored cubes on a workbench, a sliding door, a drawer, a light bulb operated with a control switch, and an LED controlled with a button. We randomly initialize these elements during data collection and evaluation. The demonstrations are either simulated using a pre-trained expert policy reuss2024multimodal or manually collected by an expert teacher. We also had the human expert specify the relevant objects and obtained ground-truth poses of these objects from the simulation environment.
@@
Figure 5: Results from our ablation study. In Explicit the system is trained on the position data of the marked objects, and in Implicit the system is trained on the masked images. CIVIL takes advantage of the human’s explicit and implicit guidance. We find that both components contribute to the overall effectiveness of CIVIL. Each policy is trained with $40$ demonstrations.
Figure 6: Section 5 results for visual imitation learning in the simulated CALVIN environment. We compare our proposed approach (CIVIL) to standard behavior cloning (BC) and baselines that use self-supervised features (BYOL), object-specific features (VIOLA and Task-VIOLA), or vision-language features (CLIP). In the Picking task we observe that CIVIL significantly outperforms all baselines in picking up the red block from different positions across the table. Our approach is particularly effective for block positions that are outside the training distribution (i.e., center of the table). This is likely because CIVIL understands what aspects of the image should influence its policy, making the system robust to background clutter or shifting positions. In the Sliding task, we find that CIVIL successfully learns to move the drawer or slider based on the state of the light bulb in just $40$ demonstrations. In contrast, baselines that use pre-trained features (BYOL and CLIP) are less precise in detecting the light signal, which reduces their success rate. This suggests that CIVIL can also learn to extract non-positional features (e.g., light bulb state) more efficiently by masking its images based on human-provided language prompts. Lastly, in the Stacking task CIVIL leverages both markers and language to extract the positions of the pink and blue blocks as well as the state of the light bulb. This enables the robot to stack the red block more successfully on either the pink or blue block, resulting in a significantly higher success rate.
@@
Demonstrations. At each timestep of a task demonstration we record an RGB image $y_{env}$ of size $200\times 200$ from a static camera that observes the entire manipulation environment, an egocentric RGB image $y_{ego}$ of size $84\times 84$ from a gripper-mounted camera, an $8$-dimensional robot state $x$, and a $7$-dimensional end-effector action $u$. The state includes $7$ joint angles of the robot arm and a Boolean gripper state. The action is a $6$-dimensional linear and angular velocity and a Boolean gripper actuation. Additionally, we obtain bounding boxes $\{\mathcal{B}\}$ for all objects in the simulation environment and explicit features $b$ in the form of $6$-dimensional Cartesian poses of relevant objects in that task. Each method uses a combination of these inputs to train the feature encoders and robot policy. For the Picking and Stacking tasks, we collect an equal amount of play data which includes images, bounding boxes, and relevant poses in randomly initialized scenarios, but it does not include robot states and expert actions.
@@
Ablations. We compare CIVIL with two ablation variants in the simulation environment to evaluate how the explicit and implicit feature modules contribute to CIVIL’s overall performance (Figure 5). In the explicit-only ablation training is restricted to the first phase, where the policy loss $\mathcal{L}_{\text{policy}}$ is optimized using unmasked image observations, and the explicit supervision loss $\mathcal{L}_{\text{explicit}}$ is applied to marker data. By contrast, the implicit-only ablation trains the causal network without incorporating $\mathcal{L}_{\text{explicit}}$ during the first training phase. For each task, the policies are trained using $40$ demonstrations. Observations from the rollouts suggest that the visual markers are particularly useful for recognizing and localizing the target object when it is not yet directly visible in the gripper camera. Policies trained with $\mathcal{L}_{\text{explicit}}$ tend to remain closer to the desired trajectory during the early stages of a rollout. However, in tasks involving multiple relevant objects such as stacking, the explicit features require more diverse play data to generalize effectively. The implicit features are designed to capture color or $2$-dimensional segmentation patterns to complement the information not represented in the $3$-dimensional spatial coordinates learned by explicit features. In practice, we found that implicit-only ablations utilizing causal networks learned richer representations from gripper images, enabling more precise stacking behaviors. Combining explicit and implicit features helps CIVIL achieves a highest success rate across all three tasks.
Results. Our results are summarized in Figure 6. Each method is trained on datasets having $10$, 2020, $40$, $80$, and $120$ demonstrations. For statistical robustness, we conduct $10$ independent training and testing runs for each method and dataset size. In each run, we test the trained policy across $100$ randomized task configurations and report the average success rate.
@@
Surprisingly, we found that standard behavior cloning performed well in this task. We attribute this result to the small size of its feature space. While BC only extracts one feature token for each camera, the object-centric methods extract two tokens: global and object-specific features. Following our analysis in Section 3.1, a more compact feature space could enable BC to learn more efficiently. Overall, this simulation result illustrates that by masking images based on language prompts CIVIL is able to extract non-positional features that help it perform the task more successfully.